# MINDS: Detection of an inner gas disk caused by evaporating bodies around HD 172555

### v2

@@ Abstract
Mechanisms such as collisions of rocky bodies or cometary activity give rise to dusty debris disks. Debris disks trace the leftover building blocks of planets, and thus also planetary composition. HD 172555, a stellar twin of Pic, hosts a debris disk thought to have resulted from a giant collision. It is known for its extreme mid-infrared silica dust feature, indicating a warm population of silica-rich grains in the asteroid belt (5 au), cold CO observed by ALMA, and small bodies evaporating as they approach close to the star. Our JWST MIRI/MRS observations now reveal emission from an inner gaseous disk (0.5 au) that arises from the evaporation of close-in material. For the first time in a debris disk, we detect neutral atomic chlorine and sulfur, as well as ionized nickel. We recovered the neutral sulfur line in 20-year-old Spitzer data, showing it is long-lived and stable. Ionized iron, previously seen only in Pic, is also detected. All lines are broadened by Keplerian rotation, pinpointing the gas location. The HD 172555 system serves as a unique laboratory to study the composition of planetesimals, asteroids, and comets beyond the Solar System. The comparison to Pic reveals, that the gas in HD 172555 is hotter, closer to the star, and poor in argon – suggesting it originates from evaporating rocky bodies near the star, while Pic’s gas may trace volatile-rich bodies from larger separations.
@@ 1 Introduction
Stars that have undergone planet formation can host detectable debris disks. The dust in these disks is primarily produced either by a continuous collisional cascades, or a singular massive collision taking place in the disk, however, cometary activity can also produce circumstellar material (e.g., Wyatt, 2008; Hughes et al., 2018). Debris disks are most often identified by excess infrared emission arising from dust, and little to no gas. They can have warm and cold components, where the temperature correlates with their separation from the star. In the solar system these correspond to the main asteroid belt at 2–3.5 au and the Edgeworth-Kuiper belt between 30 and 48 au (Wyatt, 2008). Interferometric observations sometimes also show evidence of hot exozodii, namely hot dust very close to the star (Kral et al., 2017). Infrared excess at 22–24 m around main-sequence stars is relatively common – detectable around 25% of main-sequence stars, up to 50% depending on the sample (Chen et al., 2020). However, warm dust asteroid belts that show detectable strong solid-state dust emission features (so called 10 m features) are rarer (e.g., Chen et al., 2007; Mittal et al., 2015). These features originate from large amounts of small micrometer-sized grains at about 200–300 K (Lisse et al., 2009; Johnson et al., 2012) and are thought to be a potential result of recent massive collisions or a giant impact occurring in these systems.
In recent years it has been discovered that a subset of debris disks also contains volatiles in the form of cold CO gas (e.g., Kóspál et al., 2013; Moór et al., 2017; Hughes et al., 2018; Moór et al., 2019). There is still debate about its origin as either primordial or a byproduct of the collisions in the debris disk (e.g., Smirnov-Pinchukov et al., 2022a; Cataldi et al., 2023).
The focus of our study is HD 172555 (catalog ), a nearby ( pc, Gaia Collaboration, 2020) A7V (Gray et al., 2006) star and million year old member of the Pic moving group (e.g., Barrado y Navascués et al., 1999; Miret-Roig et al., 2020). It is one of the best studied warm belt systems and a near stellar twin to Pic (A6V, Gray et al., 2006). Both stars host close to edge-on debris disks (HD 172555: 76.5∘ (Engler et al., 2018), Pic: 89∘ (Heap et al., 2000)).
@@ 1.1 Dust
HD 172555 has a strong excess in the mid-IR (; Mittal et al., 2015), typical of young A-type stars (Su et al., 2006). The fractional IR excess is a factor of 3–4 lower than in Pic (). However, while HD 172555 primarily has a strong warm dust component (about 200–350 K) corresponding to an asteroid-belt analogue, it lacks a significant cold and extended dust component (100 K). Pic hosts both warm and cold dust components, but the spectral energy distribution (SED) peaks at 60 m (Chen et al., 2007) and is dominated by the cold dust component. The warm dust of HD 172555 has been spatially resolved in the mid-infrared with TReCS at Gemini-South (Smith et al., 2012) and in the visible with polarimetric imaging using SPHERE/ZIMPOL at the VLT (Engler et al., 2018). These studies show that the dust outer radius is between 8.5 and 11.3 au (Engler et al., 2018). Based on the SED and dust models, Lisse et al. (2009) expect the bulk of the warm dust to be located at around au. A dust mass for sub-millimeter dust of about – kg, equivalent to a 150–200 km radius asteroid, was used to explain the mid-IR spectrum (Lisse et al., 2009). The millimeter dust mass was estimated to be (Schneiderman et al., 2021). HD 172555 has strongly pronounced silica and silicate features that shows an unusual composition indicative of minerals that are created at high temperatures, consistent with obsidian and tektite (Lisse et al., 2009; Johnson et al., 2012). This led to the hypothesis that the debris disk could be the direct result of a recent hypervelocity impact between two large bodies, as relative velocities of less than 10 km s-1 would not produce the observed silica (SiO2), but rather the more typical magnesium (and/or iron) containing silicates olivine and pyroxene. Such a high-velocity impact would completely melt the incident body and strip the mantel material off the impacted object (Lisse et al., 2009).
The dust features detected in the HD 172555 disk differ from the Pic disk spectrum, which is dominated by magnesium silicate features (Chen et al., 2024). These features, as noted by Chen et al. (2024), are observed to weaken over time, indicating creation of dust in giant collisions and subsequent dispersal by radiative blowout. On the other hand, the HD 172555 dust features are observed to be very stable over at least two decades (Su et al., 2020). This could be explained by a combination of mineralogy – silica particles are more transparent in the near-infrared, making them less susceptible to radiation pressure (e.g., Artymowicz, 1988; Johnson et al., 2012) – and the presence of large amounts of sub-m-grains, small enough that their optical properties change again to make them resistant to blow-out (Su et al., 2020).
@@ 1.2 Exocomet activity
High resolution spectroscopy in the optical and UV has shown non-photospheric absorption features with a stable as well as variable component indicating the presence of released gas from star-grazing evaporating bodies, also called “falling evaporating bodies” (FEB), in the HD 172555 system. In the literature the terms exocomets and FEBs are often used interchangeably, regardless of whether the evaporating bodies are predominantly icy or rocky in nature. In this work we will differentiate between falling evaporating comets and asteroids, based on their supposed reservoir of origin.
In HD 172555, Kiefer et al. (2014) have established the presence of stable and variable [Ca ii] and [Na ii] absorption at radial velocities consistent with the star. Grady et al. (2018) detected absorption lines of [Si iii] and [Si iv], [C ii] and [C iv], and [O i] using the HST/COS and STIS instruments that they associate with star-grazing objects with high covering factors up to 58%–68% of the stellar light. Furthermore, Kiefer et al. (2023) found hints of an exocomet transit in CHEOPS lightcurves, that points to a 2.5 km radius evaporating body passing at R⊙ (or au) close to the star. These detections are reminiscent of the well-studied circumstellar absorption in the Pictoris system, which has been monitored for more than 25 years (e.g. Vrignaud et al., 2024, and references therein). Transiting exocomets have also been observed in Pic (Zieba et al., 2019), underscoring a broader phenomenon of evaporating bodies around young stars. Given the extensive observational history of Pic, it provides an invaluable baseline for comparing system architectures.
@@ 1.3 Gas
CO gas has been observed by ALMA in HD 172555 in a ring from 4.2–10.8 au (Schneiderman et al., 2021) and was taken as further evidence for a recent giant impact between volatile-rich bodies, as asteroids at these separations should not retain CO. Schneiderman et al. (2021) determine a CO mass of M⊕ for gas temperatures between 100 K and 250 K. An optically thick cold disk model can result in much higher CO masses of M⊕, but seems to be ruled out from new high-resolution ALMA data (Matra et al, in prep). However, the newer data also indicate that, while the bulk of CO gas is located at a similar separation to the warm dust belt, the CO disk could extend closer to the star. No other molecules have been found using ALMA in this disk (Smirnov-Pinchukov et al., 2022b).
@@
In this paper we present the first JWST MIRI MRS observations of HD 172555. They allow us to conduct the most sensitive search for mid-infrared gas emission lines around HD 172555 to date at much higher spectral resolution. Section 2 describes the data and data reduction. Section 3 presents the emission lines we have discovered and discusses their nature and likely origin as evaporating rocky material close to the star. In Section 4 we discuss our discoveries in the context of other observations and compare the results to HD 172555’s twin star Pic. Finally, we conclude in Section 5.
@@ 2 Data reduction
We observed HD 172555 with the JWST/Mid-InfraRed Instrument (MIRI, Wright et al., 2015) in Medium Resolution Spectroscopy (MRS) mode on 18 September, 2023 as part of the MIRI Mid-Infrared Disk Survey (MINDS, PID: 1282, PI: T. Henning; Kamp et al., 2023; Henning et al., 2024). The observations made use of the four-point point-source dither pattern and has a total integration time about 10 minutes per grating. MRS provides spectral coverage from 4.9 to 27.9 m at a resolution of R1800–4000 (Jones et al., 2023; Pontoppidan et al., 2024). We processed the data using the JWST pipeline (v1.15.1; Bushouse et al., 2023) and Calibration Reference Data System context jwst_1238.pmap together with the MINDS pipeline (Christiaens et al., 2024). In the Spec2 pipeline we made use of ‘pixel_replace’ option using the minimum gradient (‘mingrad’) method to reduce the impact of bad pixels on the final spectrum. In Spec3 the spectrum is extracted band-by-band using ‘extract1d’ with automatic centroiding and spectrum level residual fringe correction enabled. The Spec3 pipeline extracts a spectrum by summing the signal in an aperture centered on the source and with a radius of 2 the full width at half maximum ( with m) of the point-spread function (PSF), performs annulus background subtraction, and aperture correction. We masked out the regions around significantly detected lines in the Spec2 residual fringe correction, which had a noticeable impact on the line amplitudes.
@@ 2.1 Spitzer IRS data reduction
HD 172555 was observed using the InfraRed Spectrograph (IRS) onboard the Spitzer Space Telescope on March 22, 2004 (AOR 3563264) and on November 6, 2007 (AOR 24368384), see Su et al. (2020) for a detailed description of the data. Spectroscopic data were obtained using the Short-Low (SL) module in all spectral orders (5.2–14 m), the second order of the Long-Low (LL) module (14–21.3 m), as well as the Short-High (SH) module (10–20 m) and the Long-High (LH) module (20–35 m).
@@
We estimate the accuracy of our flux estimates to be at the level of a few %, consistent with the uncertainties on the used stellar models of the calibrations stars. After these steps, we noticed a 2% flux difference between the short wavelength modules and the long-wavelength modules of the low-resolution spectrograph which we corrected by shifting the long-wavelength spectra downwards by this amount.
3 Results & Analysis
The new JWST/MIRI MRS spectrum and re-analyzed Spitzer IRS spectrum of HD 172555 are presented in Fig. 1. Fine-structure lines with significant detections are labeled accordingly. The combined blackbody of the stellar photosphere and warm dust components is shown as a dashed line. The pink line shows the Spitzer IRS low-resolution spectrum (see Sect. 2.1). The zoom-ins on [Ni ii], [Cl i], [Fe ii], and [S i] show that the atomic lines are broadened or even double peaked at the higher velocity resolution below 20 m. The continuum subtracted spectrum is shown in the Appendix (Fig. 5), compared to a continuum subtracted spectrum Pictoris (see Worthen et al., 2024) scaled to the distance of HD 172555 to highlight differences. The line fitting procedure and continuum subtraction is described in more detail in the Appendix B.
The JWST MIRI/MRS spectrum allows us to search for mid-infrared line emission in debris disks with much greater sensitivity and spectral resolution than Spitzer. For the first time we detect atomic line emission in the mid-IR spectrum of a debris disk other than argon and iron, which were recently observed in Pic (Worthen et al., 2024; Wu et al., 2025).
@@ 3.2 Keplerian broadening
All lines detected show significant broadening compared to the instrument resolution. The instrument resolution was measured with water emission and CO absorption lines for the whole wavelength range of MIRI MRS (Pontoppidan et al., 2024; Banzatti et al., 2025), as well as measurements for the same line transitions in protostellar atomic outflows (see Appendix Sect. B.3 and Table 3). The lines in channels 1 to 3 (18 m) show a profile that requires two components to fit, whereas lines in channel 4 show a single-peaked but highly broadened profile as expected due to the lower spectral resolution in this MRS channel.
@@
The center of the measured line profiles do not appear to be significantly shifted with respect to the stellar velocity (Table 2). We therefore rule out interstellar or background contamination as their source. This result also indicates that they are likely stable, as transient events may favor radial velocities consistent with a single offset velocity component. This behavior is similar to some of the stable lines detected in the Pic system, which show only Keplerian rotation without offset (Brandeker et al., 2004). The line profiles are consistent with Keplerian profiles that indicate a close-in gas disk generally with inner radii of 0.2 au and outer radii 0.3 au, except for [S i]. The [Ni ii], [Fe ii], and [Cl i] lines all indicate similar gas dynamics, whereas [S i] appears to be coming from larger separations (0.5 au) in addition to appearing slightly asymmetric in the line profile (see Sect. 3.3.3).
@@
Note. — [Fe ii] at 26 m is too low signal-to-noise to derive parameters reliably and is therefore not included. The values and uncertainties corresponds to the median and 16th–84th percentile ranges of the posterior. Upper limits correspond to the 90% credible interval of the posterior. a) Best-fit radial velocity of the Keplerian profile with respect to the star. No significant overall shift is detected.

### v3

@@ Abstract
Mechanisms such as collisions of rocky bodies or cometary activity give rise to dusty debris disks. Debris disks trace the leftover building blocks of planets, and thus also planetary composition. HD 172555, a stellar twin of Pic, hosts a debris disk thought to have resulted from a giant collision. It is known for its extreme mid-infrared silica dust feature, indicating a warm population of silica-rich grains in the asteroid belt (5 au), cold CO observed by ALMA, and small bodies evaporating as they approach close to the star. Our JWST MIRI MRS observations now reveal emission from an inner gaseous disk (0.5 au) that arises from the evaporation of close-in material. For the first time in a debris disk, we detect neutral atomic chlorine and sulfur, as well as ionized nickel. We recovered the neutral sulfur line in 20-year-old Spitzer data, showing it is long-lived and stable. Ionized iron, previously seen only in Pic, is also detected. All lines are broadened by Keplerian rotation, pinpointing the gas location. The HD 172555 system serves as a unique laboratory to study the composition of planetesimals, asteroids, and comets beyond the Solar System. The comparison to Pic reveals that the gas in HD 172555 is hotter, closer to the star, and poor in argon – suggesting it originates from evaporating rocky bodies near the star, while Pic’s gas may trace volatile-rich bodies from larger separations.
@@ 1 Introduction
Stars that have undergone planet formation can host detectable debris disks. The dust in these disks is primarily produced either by a continuous collisional cascade, or a singular massive collision taking place in the disk; however, cometary activity can also produce circumstellar material (e.g., Wyatt, 2008; Hughes et al., 2018). Debris disks are most often identified by excess infrared emission arising from dust, and little to no gas. They can have warm and cold components, where the temperature correlates with their separation from the star. In the Solar System, these correspond to the main asteroid belt at 2–3.5 au and the Edgeworth–Kuiper belt between 30 and 48 au (Wyatt, 2008). Interferometric observations sometimes also show evidence of hot exozodii, namely hot dust very close to the star (Kral et al., 2017). Infrared excess at 22–24 m around main-sequence stars is relatively common – detectable around 25% of main-sequence stars, up to 50% depending on the sample (Chen et al., 2020). However, warm dust asteroid belts that show detectable strong solid-state dust emission features (so-called 10 m features) are rarer (e.g., Chen et al., 2007; Mittal et al., 2015). These features originate from large amounts of small micrometer-sized grains at about 200–300 K (Lisse et al., 2009; Johnson et al., 2012) and are thought to be a potential result of recent massive collisions or a giant impact occurring in these systems.
In recent years, it has been discovered that a subset of debris disks also contains volatiles in the form of cold CO gas (e.g., Kóspál et al., 2013; Moór et al., 2017; Hughes et al., 2018; Moór et al., 2019). There is still debate about its origin as either primordial or a byproduct of the collisions in the debris disk (e.g., Smirnov-Pinchukov et al., 2022; Cataldi et al., 2023).
The focus of our study is HD 172555 (catalog ), a nearby ( pc, Gaia Collaboration, 2020) A7V (Gray et al., 2006) star and million year old member of the Pic moving group (e.g., Barrado y Navascués et al., 1999; Miret-Roig et al., 2020). It is one of the best studied warm belt systems and a near stellar twin to Pic (A6V; Gray et al., 2006). Both stars host close to edge-on debris disks (HD 172555: 76.5∘ (Engler et al., 2018), Pic: 89∘ (Heap et al., 2000)).
@@ 1.1 Dust
HD 172555 has a strong excess in the mid-IR (; Mittal et al., 2015), typical of young A-type stars (Su et al., 2006). The fractional IR excess is a factor of 3–4 lower than in Pic (). However, while HD 172555 primarily has a strong warm dust component (about 200–350 K) corresponding to an asteroid-belt analog, it lacks a significant cold and extended dust component (100 K). Pic hosts both warm and cold dust components, but the spectral energy distribution (SED) peaks at 60 m (Chen et al., 2007) and is dominated by the cold dust component. The warm dust of HD 172555 has been spatially resolved in the mid-IR with TReCS at Gemini-South (Smith et al., 2012) and in the visible with polarimetric imaging using SPHERE/ZIMPOL at the VLT (Engler et al., 2018). These studies show that the dust outer radius is between 8.5 and 11.3 au (Engler et al., 2018). Based on the SED and dust models, Lisse et al. (2009) expected the bulk of the warm dust to be located at around au. A dust mass for submillimeter dust of about – kg, equivalent to a 150–200 km radius asteroid, was used to explain the mid-IR spectrum (Lisse et al., 2009). The millimeter dust mass was estimated to be (Schneiderman et al., 2021). HD 172555 has strongly pronounced silica and silicate features that show an unusual composition indicative of minerals that are created at high temperatures, consistent with obsidian and tektite (Lisse et al., 2009; Johnson et al., 2012). This led to the hypothesis that the debris disk could be the direct result of a recent hypervelocity impact between two large bodies, as relative velocities of 10 km s-1 would not produce the observed silica (SiO2), but rather the more typical magnesium (and/or iron) containing silicates olivine and pyroxene. Such a high-velocity impact would completely melt the incident body and strip the mantel material off the impacted object (Lisse et al., 2009).
The dust features detected in the HD 172555 disk differ from the Pic disk spectrum, which is dominated by magnesium silicate features (Chen et al., 2024). These features, as noted by Chen et al. (2024), are observed to weaken over time, indicating creation of dust in giant collisions and subsequent dispersal by radiative blowout. On the other hand, the HD 172555 dust features are observed to be very stable over at least two decades (Su et al., 2020). This could be explained by a combination of mineralogy – silica particles are more transparent in the near-infrared, making them less susceptible to radiation pressure (e.g., Artymowicz, 1988; Johnson et al., 2012) – and the presence of large amounts of submicron grains, small enough that their optical properties change again to make them resistant to blowout (Su et al., 2020).
@@ 1.2 Exocomet activity
High-resolution spectroscopy in the optical and UV has shown nonphotospheric absorption features with a stable as well as variable component indicating the presence of released gas from star-grazing evaporating bodies, also called “falling evaporating bodies” (FEBs), in the HD 172555 system. In the literature, the terms exocomets and FEBs are often used interchangeably, regardless of whether the evaporating bodies are predominantly icy or rocky in nature. In this work, we will differentiate between falling evaporating comets and asteroids, based on their supposed reservoir of origin.
In HD 172555, Kiefer et al. (2014) established the presence of stable and variable [Ca ii] and [Na ii] absorption at radial velocities consistent with the star. Grady et al. (2018) detected absorption lines of [Si iii] and [Si iv], [C ii] and [C iv], and [O i] using the Hubble Space Telescope (HST)/Cosmic Origins Spectrograph (COS) and STIS instruments that they associate with star-grazing objects with high covering factors up to 58%–68% of the stellar light. Furthermore, Kiefer et al. (2023) found hints of an exocomet transit in CHEOPS lightcurves, that points to a 2.5 km radius evaporating body passing at R⊙ (or au) close to the star. These detections are reminiscent of the well-studied circumstellar absorption in the Pictoris system, which has been monitored for more than 25 yr (e.g., Vrignaud et al., 2024, and references therein). Transiting exocomets have also been observed in Pic (Zieba et al., 2019), underscoring a broader phenomenon of evaporating bodies around young stars. Given the extensive observational history of Pic, it provides an invaluable baseline for comparing system architectures.
@@ 1.3 Gas
CO gas has been observed by the Atacama Large Millimeter/submillimeter Array (ALMA) in HD 172555 in a ring from 4.2–10.8 au (Schneiderman et al., 2021) and was taken as further evidence for a recent giant impact between volatile-rich bodies, as asteroids at these separations should not retain CO. Schneiderman et al. (2021) determined a CO mass of M⊕ for gas temperatures between 100 and 250 K. An optically thick cold disk model can result in much higher CO masses of M⊕. New higher spatial resolution ALMA data indicate that, while the bulk of CO gas is located at a similar separation to the warm dust belt, the CO disk could extend closer to the star (L. Matra, private communication). No other molecules have been found using ALMA in this disk (Smirnov-Pinchukov et al., 2022).
@@
In this paper, we present the first JWST Mid-InfraRed Instrument (MIRI) Medium Resolution Spectroscopy (MRS) observations of HD 172555. They allow us to conduct the most sensitive search for mid-IR gas emission lines around HD 172555 to date at much higher spectral resolution. Section 2 describes the data and data reduction. Section 3 presents the emission lines we have discovered and discusses their nature and likely origin as evaporating rocky material close to the star. In Section 4 we discuss our discoveries in the context of other observations and compare the results to HD 172555’s twin star Pic. Finally, we conclude in Section 5.
@@ 2 Data reduction
We observed HD 172555 with the JWST/MIRI (Wright et al., 2015) in MRS mode on 2023 September 18 as part of the MIRI Mid-Infrared Disk Survey (MINDS, PID: 1282, PI: T. Henning; Kamp et al., 2023; Henning et al., 2024). The observations made use of the four-point point-source dither pattern and have a total integration time of about 10 minutes per grating. MRS provides spectral coverage from 4.9–27.9 m at a resolution of R1800–4000 (Jones et al., 2023; Pontoppidan et al., 2024). We processed the data using the JWST pipeline (v1.15.1; Bushouse et al., 2023) and Calibration Reference Data System context jwst_1238.pmap together with the MINDS pipeline (Christiaens et al., 2024). In the Spec2 pipeline, we made use of ‘pixel_replace’ option using the minimum gradient (‘mingrad’) method to reduce the impact of bad pixels on the final spectrum. In Spec3 the spectrum is extracted band-by-band using ‘extract1d’ with automatic centroiding and spectrum level residual fringe correction enabled. The Spec3 pipeline extracts a spectrum by summing the signal in an aperture centered on the source and with a radius of 2 the FWHM of the point-spread function (PSF), performs annulus background subtraction, and performs aperture correction. We masked out the regions around significantly detected lines in the Spec2 residual fringe correction, which had a noticeable impact on the line amplitudes.
@@ 2.1 Spitzer IRS data reduction
HD 172555 was observed using the InfraRed Spectrograph (IRS) on board the Spitzer Space Telescope on 2004 March 22 (AOR 3563264) and on 2007 November 6 (AOR 24368384), see Su et al. (2020) for a detailed description of the data. Spectroscopic data were obtained using the Short-Low module in all spectral orders (5.2–14 m), the second order of the Long-Low module (14–21.3 m), as well as the Short-High module (10–20 m) and the Long-High module (20–35 m).
@@
We estimate the accuracy of our flux estimates to be at the level of a few percent, consistent with the uncertainties on the used stellar models of the calibrations stars. After these steps, we noticed a 2% flux difference between the short-wavelength modules and the long-wavelength modules of the low-resolution spectrograph, which we corrected by shifting the long-wavelength spectra downward by this amount.
3 Results and Analysis
The new JWST/MIRI MRS spectrum and reanalyzed Spitzer IRS spectrum of HD 172555 are presented in Fig. 1. Fine-structure lines with significant detections are labeled accordingly. The combined blackbody of the stellar photosphere and warm dust components is shown as a dashed line. The pink line shows the Spitzer IRS low-resolution spectrum (see Sect. 2.1). The zoom-ins on [Ni ii], [Cl i], [Fe ii], and [S i] show that the atomic lines are broadened or even double peaked at the higher velocity resolution below 20 m. The continuum-subtracted spectrum is shown in Appendix E (Fig. 5), compared to a continuum-subtracted spectrum of Pictoris (see Worthen et al., 2024) scaled to the distance of HD 172555 to highlight differences. The line fitting procedure and continuum subtraction is described in more detail in Appendix B.
The JWST MIRI MRS spectrum allows us to search for mid-IR line emission in debris disks with much greater sensitivity and spectral resolution than Spitzer. For the first time, we detect atomic line emission in the mid-IR spectrum of a debris disk other than argon and iron, which were recently observed in Pic (Worthen et al., 2024; Wu et al., 2025).
@@ 3.2 Keplerian broadening
All lines detected show significant broadening compared to the instrument resolution. The instrument resolution was measured with water emission and CO absorption lines for the whole wavelength range of MIRI MRS (Pontoppidan et al., 2024; Banzatti et al., 2025), as well as measurements for the same line transitions in protostellar atomic outflows (see Appendix Sect. B.3 and Table 3). The lines in channels 1–3 (18 m) show a profile that requires two components to fit, whereas lines in channel 4 show a single-peaked but highly broadened profile as expected due to the lower spectral resolution in this MRS channel.
@@
The center of the measured line profiles does not appear to be significantly shifted with respect to the stellar velocity (Table 2). We therefore rule out interstellar or background contamination as their source. This result also indicates that they are likely stable, as transient events may favor radial velocities consistent with a single offset velocity component. This behavior is similar to some of the stable lines detected in the Pic system, which show only Keplerian rotation without offset (Brandeker et al., 2004). The line profiles are consistent with Keplerian profiles that indicate a close-in gas disk generally with inner radii of 0.2 au and outer radii 0.3 au, except for [S i]. The [Ni ii], [Fe ii], and [Cl i] lines all indicate similar gas dynamics, whereas [S i] appears to be coming from larger separations (0.5 au) in addition to appearing slightly asymmetric in the line profile (see Sect. 3.3.3).
@@
Note. — [Fe ii] at 26 m is too low of a signal-to-noise to derive parameters reliably and is therefore not included. The values and uncertainties correspond to the median and 16th–84th percentile ranges of the posterior. Upper limits correspond to the 90% credible interval of the posterior. a) Best fit radial velocity of the Keplerian profile with respect to the star. No significant overall shift is detected.